- Press Release
- Dec 8, 2022
NASA Space Station On-Orbit Status 25 Mar 2004
All ISS systems continue to function nominally, except those noted previously or below.
The Elektron oxygen generator continues to run smoothly, ever since the replacement of its BZh Liquid Unit last week. ppO2 (oxygen partial pressure) in the cabin has leveled out satisfactorily, reading 150 mmHg (Torr) this morning.
CDR/SO Michael Foale continued payload/science activities in the Lab, starting out with the PFMI (Pore Formation & Mobility Investigation) in the MSG (Microgravity Science Glovebox). [The task involved removal of the PFMI-14 sample from the thermal chamber, followed by the exchange of the PFMI laptop’s g-LIMIT (Glovebox Integrated Microgravity Isolation Technology) bypass micro drive assembly and removal of the associated videotape. Mike was commended for restoring the experiment to full operation on 3/23, jerry-rigging a low-current (<10 mA) cable connector with Kapton tape after it broke during troubleshooting of a failed hardware startup.]
At the Lab’s Maintenance Work Area (MWA), the Science Officer performed the 4th
and final isothermal experiment of the MFMG (Miscible Fluids in Microgravity) research. [Today, Foale prepared a syringe of 100% honey, into which he then pulled tinted water with the plunger. Per his request, time has been reserved tomorrow for a debrief telecon with the Principal Investigator.]
Later, Foale started out on the new BCAT-3 (Binary Colloidal Alloy Test-3) experiment for NASA GRC (Glenn Research Center)’s microgravity research program. Today’s activities focused on familiarization with uplinked crew training material, a teleconference with the BCAT team, and setting up the hardware. [BCAT-3 is a Small Payload for ISS using the Kodak DCS760 digital camera with MagLite at the MWA. Experiment hardware for homogenizing samples in micro-G include a Slow Growth Sample Module with sample couvettes and the BCAT Magnet for homogenizing the alloy samples (toxicity level 1) to initiate growth of colloidal structures. Forerunners were the glovebox investigations BCAT & BCAT-2 launched on STS-79 & STS-86 to the Russian space station Mir during the fall of ‘96 and ’97. BCAT-3 is a precursor for the LMM (Light Microscopy Module) scheduled to fly in 2006. BCAT-3 is also a follow-on experiment to CGEL (Colloidal Gelation) operated by Mike Foale on Mir/Increment 5. Possible future applications of the colloidal alloy experiments are photonic crystals for telecommunications and computer applications (e.g., optical switches and waveguides, “computing with light”), extremely low threshold lasers, and improved use of supercritical fluids (e.g. CO2 for food extractions, pharmaceuticals, dry cleaning, etc.)]
FE Alex Kaleri completed the regular periodic download of data and imagery collected of the BIO-5 Rasteniya-2 (“Plants-2”) experiment to the computer for subsequent downlink to the ground. [Rasteniya studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.]
Continuing his maintenance work on the Russian onboard television system (TVS), Kaleri replaced the #2 video monitor (VKU-2), found flawed on 3/23, with a new unit (VKU-1) from spares. Later, he performed an installation & performance test of VKU-1 (MTs-27) with ground support via S-band. [On TsUP Go, he was then to return the replaced monitor to stowage in the FGB.]
For tomorrow’s EVA/spacesuit midterm maintenance activities, Mike Foale printed out the 33-page list of procedural instructions for review. Later, the crew conducted a tagup with EVA experts on console to discuss the EMU (extravehicular mobility unit) procedures. [The mandatory midterm checkout verifies that the components of EMU 3011 are operating nominally following an extended period of downtime. It also satisfies the maintenance requirements for extended on-orbit use. The C/O includes cycling the valves and regulators in the backpack, which is performed once per year. The EMU fan needs to run for two hours, but the crew is not required to monitor the system during the run time. A PCS (portable computer system) laptop will be connected to the EMU during the C/O, in order to gather more data than can be seen in real-time on the ground. The data will later be transferred to an SSC (station support computer), to be retrieved by the ground. Prior to the C/O, the crew will remove stowage goods from the U.S. Airlock (A/L) to make room for the activities, and afterwards the A/L will be restowed. The CCAA (common cabin air assembly) air conditioner will be turned on for the A/L activities, but a subsequent heat exchanger dryout will not be required since the LTL (low temperature loop)’s temperature setpoint will be chosen high enough to prevent condensation.]
The CDR conducted a periodic 15-min. inspection of the food warmers in the galley (located in the SM). [Food warmers, food trays, utensils, etc., are part of the food preparation hardware of the Russian-provided CSS (crew support systems). The food warmers are in recessed wells in the galley table that also has crew and equipment restraints (bungees and Velcro). Besides the galley, the wardroom area also includes a potable water dispenser (for hot and ambient water for drink and food hydration), and a trash container.]
Sasha Kaleri performed the regular once-a-week maintenance reboot on the operational PCS laptops and also restarted the OCA comm router laptop (every two weeks).
The crew worked out on TVIS, RED exerciser and VELO cycle with load trainer. [As reported yesterday, the forward RED canister (#1011) has started “scraping” at low loads and exhibits more resistance than usual. At higher loads the pull of the “rowing” machine’s Flexpack canister is smooth and noiseless. #1011 & #1012 are the only two canisters on ISS that have full load capacity remaining, and the intent was to use these cans through the end of Expedition 8 (April 29). On 3/27 (Saturday), the crew will perform a 30-min. calibration check on #1011 (by pulling the cord three times at each of 7 specific load settings) and return to modified exercise protocol, which uses #1012 for all low-load exercises.]
As every week, Mike transferred data files from the physical exercise equipment to the MEC (medical equipment computer) via memory card and RED log entries, for downlink on OCA comm. Later, he completed the periodic transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, then erasing them on the HRM.
The FE meanwhile took the periodic reading of the cabin air’s current carbon dioxide partial pressure in the SM and Lab, using the U.S. CDMK (CO2 monitor kit), for calldown to MCC-Houston (along with the battery status) for use in trending analyses.
Sasha also prepared the daily IMS inventory “delta” file for automated updating the IMS databases. The scheduled daily inspection/maintenance of the SM’s SOZh life support system (including ASU toilet facilities) was deferred.
At 7:35am EST, Foale and Kaleri conducted a PAO downlink via S-band and Ku-band/NetMeeting, by crew request, to send greetings to the initiation/launch of Regional Science Centers in the United Kingdom (UK).
The new permanent procedure for attitude control handovers from the RS (Russian segment) motion control system directly into a U.S. Momentum Management controller will be used for the first time on 3/26 (Friday) night, 9:23pm EST, when the ISS maneuvers from XPOP (x-axis perpendicular to orbit plane) back to LVLH XVV (local vertical local horizontal/x-axis in velocity vector). [The automated procedure allows going from RS thruster control straight to U.S. CMG momentum management without the previously required period of attitude hold. The change is intended to reduce the stress on the CMGs (control moment gyros) during activities calling for the handover, and it will normally require no crew support (except perhaps when Russian command steps are to be performed in the absence of TsUP).]
Since Progress 13P hatch closing is scheduled quite soon after arrival of Expedition 9 (4/18 EDT), the crew was advised to start pre-gathering and temporarily stowing trash items in mesh bags and/or CTBs (crew transfer bags) for later stowage on the cargo drone turned trash container.
According to specialists (who met on 3/22), the data indicating an air leak into the Lab science window ‘s inter-pane “Volume D” is believed to be correct. It still needs to be determined if the much-higher-than-expected pressure increase is due to an actual leak or if it was introduced during the evacuation on 3/5. [The current Lab window operating constraints are still considered valid for protecting the window’s optical qualities, even if the apparent leak is real.]
Kaleri’s checkout on 3/22 of the TORU teleoperator control system between the SM and the docked Progress-260/13P (to help analyze problems reported with TORU during 13P docking on 1/3) revealed that the primary Russian VHF (very-high frequency) receiver indeed is not functioning. The backup receiver was found to be OK.
During inspection of recently acquired photos of the portside forward PVR (photovoltaic radiator)’s base-plate for a known “bubble” more surface irregularities near the bubble and on the starboard side base plate were discovered. A photo survey of the port base plate near the bubble is already planned for 4/1, and a survey of the starboard side is under discussion.
Today’s CEO (Crew Earth Observations) targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were Tropical Cyclone Fay (just off NW Australia, was worthy of an image. Dynamic events have a short shelf life, so that rapid downlink is ideal for possible posting to the web), Cyclone Oscar, Indian Ocean (Dynamic event. Nadir pass over this new storm that will become a cyclone), Singapore (nadir pass over the island nation at the tip of the Malay Peninsula), Internal waves, Vietnam (looking left into the glint point for internal waves developed along the bulge of Vietnam), Betsiboka River Delta, Madagascar (good near-nadir pass over this largest estuary on Madagascar, in the aftermath of floods by a tropical cyclone), Bombay Region, India (looking right towards the coast of India to document reported aerosols. Visibility in Bombay is reduced to 2 miles), Ganges River Basin (looking left and right for an aerosol blanket. 2-mile visibility also reported in cities on the Ganges River), Omo River Delta, Ethiopia (400mm-lens. Nadir pass over the delta that occupies the north end of Lake Turkana. Changes in delta size are a prime proxy for long-term rainfall in Ethiopia, the Nile headwaters), Addis Ababa, Ethiopia (looking just left of track for this remote capital city), Amman, Jordan (nadir pass over the Jordanian capital), Patagonian Glaciers (opportunity between cloud masses for views of smaller glaciers on the east, cloudfree flanks of the southern Andes), and Panama Canal (400mm-lens. Pass just east of the canal. Looking left for a detailed view).
CEO images can be viewed at the websites.
See also the website “Space Station Challenge” at
U.S. and Russian Segment Status (as of today, 1:35pm EST).
Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
- Elektron O2 generator is On. Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring. MCA (major constituents analyzer) is in Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
- SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 26.3; ppO2 (mmHg) — 148.0; ppCO2 (mmHg) — 3.8;
- SM Transfer Compartment: Pressure (mmHg) — 743; temperature (deg C) — 20.2.
- FGB Cabin: Pressure (mmHg) — 744; temperature (deg C) — 23.3.
- Node: Pressure (mmHg) — 745.16; temperature (deg C) — 24.1 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- U.S. Lab: Pressure (mmHg) — 747.14; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- Joint Airlock (Equip. Lock): Pressure (mmHg) — 747.34; temperature (deg C) — 25.3; shell heater temp (deg C) — 24.2, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- PMA-1: Shell heater temp (deg C) — 24.9
- PMA-2: Shell heater temp (deg C) — 13.2
(n/a = data not available)
Electrical Power Systems (EPS):
- Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Autotrack (suntracking) and bias-angled 43 deg. for drag reduction (“sun slicer”)
- SM batteries: Battery #8 is off-line; all other batteries (7) are in “Partial Charge” mode.
- FGB batteries: Battery #6 is off-line (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
- Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.
Command & Data Handling Systems:
- C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
- GNC-1 MDM is prime; GNC-2 is Backup.
- INT-2 is operating; INT-1 is Off.
- EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
- LA-1, LA-2 and LA-3 MDMs are all operating.
- PL-1 MDM is Off; PL-2 MDM is Operational.
- APS-1 (automated payload switch #1) and APS-2 are both On.
- SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
- SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
- Total propellant load available: 3926 kg (8655 lb) as of 3/18/04; [SM(755) + FGB(2512) + Progress M-1(659)]. (Capability: SM — 860 kg; FGB — 6120 kg).
Attitude Control Systems:
- 3 CMGs on-line (CMG-1 failed).
- State vector source — U.S. SIGI-1 (GPS)
- Attitude source — U.S. SIGI-1 (GPS)
- Angular rate source — RGA-1
- XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist), until 3/28.
Communications & Tracking Systems:
- FGB MDM-1 is powered Off; FGB MDM-2 is operational.
- All other Russian communications & tracking systems are nominal.
- *S-band is operating nominally (on string 2).
- Ku-band is operating nominally.
- Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
- Video subsystem operating nominally.
- HCOR (high-rate communications outage recorder) is operating nominally.
- SSRMS/Canadarm2 based at Lab PDGF/LEE A, powered on both strings.
- MBS: KA power on both strings.
- MT: latched and mated at WS4.
- POA: KA power on both strings.
- RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.
ISS Orbit (as of this morning, 5:44am EST [= epoch]):
- Mean altitude — 366.7 km
- Apogee — 373.9km
- Perigee — 359.5 km
- Period — 91.88 min.
- Inclination (to Equator) — 51.6303 deg
- Eccentricity — 0.001065
- Solar Beta Angle — ~25 (magnitude decreasing)
- Orbits per 24-hr. day — 15.67
- Mean altitude loss last 24 hours — 140 m
- Revolutions since FGB/Zarya launch (Nov. ’98) — 30517
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see